OpenStudy (anonymous):

hwat is the value of k if the system has no solution: 2x-y=3 4x+ky=4

OpenStudy (anonymous):

2x = 3 + y 2(3+y) + ky = 4 6 + 2y + ky = 4 y(2+k) = -2 y = -2 / (2+k) k = -2

OpenStudy (anonymous):

coolsector has a good answer, but you can do it with your eyeball. 4x is twice 2x so make sure that ky is twice -y i.e. k = -2